OpenStudy (anonymous):

How would you solve this system?: u-t=5 10t+u=3(t+u)

hero (hero):

I guess you didn't understand my steps

OpenStudy (anonymous):

I just don't get the 10t+u=3(t+u) part.

hero (hero):

It's not that difficult. Think outside the box

OpenStudy (anonymous):

There's actually a difference between the other question. Can other people try and respond to this question as well?

Directrix (directrix):

u-t=5 10t+u=3(t+u) -------------------- u = 5 + t Substitution 10t + (5 + t) = 3 ( t + (5 + t) ) 10t + 5 + t = 3t + 15 + 3t 11t + 5 = 6t + 15 5t = 10 t = 2 u = 5 + t u = 5 + 2 = 7 Solution ( 2 , 7 )

OpenStudy (anonymous):

Thank you!

Directrix (directrix):

Glad to help. Thanks for the medal.